OpenStudy (anonymous):

Students use a plastic bow and arrow set in a laboratory experiment to examine projectile motion. The bowstring is pulled back a distance of 16.2cm to launch the arrow, which has a mass of 75g. When the arrow is fired horizontally from a height of 1m above the floor, it lands on the lfoor a horizontal distance of 6.8m from the bow. (a) Determine the launch velocity of the arrow. (b) caclulate the elastic potential energy when the arrow is pulled back ready to launch. (c) Determine the elastic constant,k, of the bowstring. (d)determine the speed of the arrow when it hits the floor. @M

OpenStudy (amistre64):

since we are working with meters gravity = -9.8

OpenStudy (amistre64):

-g/2 t^2 + Vi t + Hi is the equation for the amount of time it takes to hit the ground

OpenStudy (amistre64):

Hi = 1 in this case and as before g = 9.8 which simplifies to -4.9

OpenStudy (amistre64):

suppose Vi, initial velocity is 0; then calculate how long it takes to drop 1 meter: -4.9 t^2 + 1 = 0 when t^2 = 1/4.9

OpenStudy (amistre64):

or rather t= sqrt(1/4.9)

OpenStudy (amistre64):

horizontal distance is measured by (time) * speed = distance

OpenStudy (amistre64):

that might help out some

OpenStudy (amistre64):

velocity = 6.8/sqrt(1/4.9) in this case: which is abt 15.05 meters per sec.

OpenStudy (anonymous):

thanks i just calculated that, works out. i can get the rest now lol

OpenStudy (amistre64):

good, cause I got no idea about the rest of it lol